guys i have trouble in binomial theorem... all the sigma and double sigma problems.... how to solve them ?? urrgent help tests comin up !! 

Grade:11

2 Answers

AskiitiansExpert Mohit-IITD
13 years ago

Dear Aditya,

There are different ways to solve sigma based problems.

For Sigma based problems, you need to first understand what ∑ means. It is the summation of all the terms e.g for i=1 to n, ∑i=1+2+3+...+n=n(n+1)/2.

Now for binomial, there are various formulae which you must learn like ∑nCr=2n for r=0 to n etc.

Similarly for double summation, ∑∑ it is carried out over 2 variables, say 'r' and 's', so we simply carry out the inner summation first and then the outer summation. (as i or r earlier) 

I would recommend you to go through http://www.askiitians.com/iit-jee-algebra/binomial-theorem-for-a-positive-integral-index/ to understand this topic further. You will need to spend about 6-8 hours to understand the topic better.

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation. 

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

sahil
10 years ago
hey first carried out summation,then move onto..inner one....